OpenStudy (anonymous):

Need more help...If 2x-5y=8 and 7x+6y=17, then 9x+y=

OpenStudy (anonymous):

from eq 1 x =(8 + 5y)/2 from eq 2 x = (17-6y)/7

OpenStudy (radar):

Use substitution. Equation1 2x-5y=8 Solve for x, 2x=8+5y, then x=4+(5/2)y Equation 2, 7x=6y=17, substitute the newly found x expression like this; 7(4+(5/2)y)+6y=17 this becomes 28+35/2y+6y=17 consolidating similar terms 47y/2=-11 now solve for y

OpenStudy (anonymous):

I'm guessing 25

OpenStudy (radar):

After determing y substitute that value for y int equation 1 and solve for x. then solve the 3rd equation

OpenStudy (anonymous):

You can add the two equations. It works out well

OpenStudy (amistre64):

now lokis a superhero too ;)

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

Who pipped me over?

OpenStudy (amistre64):

easy, I fanned you lol

OpenStudy (anonymous):

Oh...thank you ;)

OpenStudy (amistre64):

been waiting a long time to do that lol ;)

OpenStudy (anonymous):

I reciprocated.

OpenStudy (anonymous):

I prefer the pink of hero...but not complaining.

OpenStudy (amistre64):

i here at500 we get to be villians lol

OpenStudy (amistre64):

megamind would be an awesome title ;)

OpenStudy (anonymous):

I know others answered this question, but I wrote out a solution while you were doing so, and didn't want to waste it :)

OpenStudy (anonymous):

I dunno...I wish they'd turn the fan thing off half the time.

OpenStudy (anonymous):

I think it's much easier just to add the two equations, since they sum up to: 9x+y=25

OpenStudy (anonymous):

rofl, yes!

OpenStudy (amistre64):

tag!!